OpenStudy (anonymous):

If 18 meters of high-resistance wire with a diameter of 1.2 mm has a resistance of 10 ohms, what is the resistance of 27 m of the same type of wire with a diameter of 1.5 mm?

OpenStudy (irishboy123):

\[R = constant \times \frac{ length }{ diameter^{2} }\] either calc the "constant" from the first observation and use that to calc the second resistance; or just go: \[R _{2} = R_{1} \times \frac{ d_{1}^{2} }{ l_{1} } \times \frac{ l_{2} }{ d_{2}^{2} }\]